OpenStudy (anonymous):

Find the 18th term of the sequence 5, 8, 11, 14, 17...??????????? #confused

OpenStudy (anonymous):

you have an arithmetic progression with ratio=3 and the first term 5 write the general term formula

OpenStudy (anonymous):

A18=A1+(n-1)d where d is ratio; =>A18=5+17*3=56

OpenStudy (anonymous):

The 18th term of the sequence is 56